OpenStudy (calculusxy):

Alejandro has 3 times as many $20 bills in his wallet as he has $5 bills. He also has twice as many $1 bills than $20 bills. If he has a total of $142 in his wallet, how much money does Alejandro have in $1 bills?

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

give a name to the number of $1 bills. (that is what you want to find)

OpenStudy (calculusxy):

c ?

OpenStudy (calculusxy):

is it fine to have a variable for each of the bills? $20 bills = a $5 bills = b $1 bills = c

OpenStudy (calculusxy):

I thought that I can make three equations 3a = b 2c = a a + b + c = $142

OpenStudy (phi):

yes, that is ok. But we have to "get rid" of the other 2 variables at some point. a + b + c = $142 rethink that.

OpenStudy (phi):

you have to keep it clear. a is the # of 20 dollar bills. (that is different from the # of dollars)

OpenStudy (phi):

in other words, the amount of money is the # of bills times how much each bill is worth

OpenStudy (calculusxy):

okay so how can i change the third one?

OpenStudy (phi):

think about a simpler problem and try to get that correct say you had "a" 20 dollar bills and the total dollars you have is 20 would you say a=20 ?

OpenStudy (calculusxy):

would i say it as 20a -> 20(20) ?

OpenStudy (phi):

I'm not too clear here. you have 1 $20 bill but we don't know it's 1 bill. instead we are told we have "a" bills and the total dollars is 20

OpenStudy (phi):

we would say the amount of money is 20a and that equals 20 20a= 20 (and we solve for a to get 1 , the number of bills) in your equation a+b+c= 142 a+b+c is the number of bills, but not the amount of money

OpenStudy (faiqraees):

and shouldn't the total equation be like 3x(20)+x(5)+2x(1)=142

OpenStudy (faiqraees):

oh sorry there should be 6x(1) instead of 2x(1)

OpenStudy (calculusxy):

I don't understand how to get the total equation. Can someone explain?

OpenStudy (faiqraees):

okay we're going to take ratios the ration of 20 to 5 is 3 to 1 right?

OpenStudy (phi):

if you had 1 20, 1 5 and 1 1 how much money do you have ?

OpenStudy (calculusxy):

$26

OpenStudy (phi):

yes, but a=1, b=1 and c=1 and a+b+c= 3 not 26 how do we fix that ?

OpenStudy (calculusxy):

do we have 20a + 5b + 1c = $142?

OpenStudy (phi):

yes

OpenStudy (phi):

For the other equations 3 times as many $20 bills in his wallet as he has $5 bills say he has one $5 bill, then he has 3 $20 bills b=1 a=3 it looks like a=3b (which is backwards from what you wrote up above: b=3a) these type sentences are confusing, so I try to think about it the way I just posted.

OpenStudy (faiqraees):

Since I have to leave I will type the whole method. We can approach a solution using the ratio method and from the portrayal of the question it is almost evident that they are also looking for the same method It is stated the ratio of 20 bills to 5 bill is 3 : 1 and the ratio of 20 bill to 1bill is 1:2 We can make a unified ratio by multiplying the last ratio by 3 to get a uniform ratio 20 bill to 1 bill to 5 bill 3 : 6 : 1 Now its apparent that the ratio of the bill doesn't equate to 142 but a small fraction of it we can write 3(20)+6(1)+1(5)=142/x (142/x being the fraction part that the ratio equate to. Note treat the dollar bill as variables.) Now multiply the x on both sides and voila you have the equation

OpenStudy (calculusxy):

i don't know how you got a= 3b when it was b=1 and a=3 @phi

OpenStudy (phi):

it's either a=3b or b=3a. if we know b is 1 and a is 3 only the first equation works

OpenStudy (faiqraees):